OpenStudy (anonymous):

f(x)= [2x,x<1 [5,x=1 [x^2, x>1 which of the following are true? A) f(5)=1 B) f(2) =4 C) f(-2)=4 f(1)=5

OpenStudy (anonymous):

B & D

OpenStudy (shamim):

If x=5 5>1, so x>1 f(x)=x^2 f(5)=5^2=25 So A is not correct

OpenStudy (shamim):

When x=2 2>1, so x>1 f(x)=x^2 f(2)=2^2=4 So B is correct

OpenStudy (shamim):

When x=-2 -2<1 x<1 So f(x)=2x f(-2)=2*(-2)=-4 So C is not correct

OpenStudy (shamim):

Try yourself the rest 1

OpenStudy (shamim):

Response plz